Find the eigenvalues and eigenfunctions of

𝑦

β€²β€² +



𝑦 = 0 , 0 < π‘₯ < 3

𝑦′(0) = 0 , 𝑦′(3) = 0


1
Expert's answer
2020-10-27T19:18:52-0400
"y''+\lambda y=0, y'(0)=0,y'(3)=0"


The auxiliary equation


"r^2 +\lambda=0"

"r=\pm\sqrt{-\lambda}, \lambda\leq0"

The corresponding solution to our ODE will be


"y=A\cos(\sqrt{\lambda }x)+B\sin(\sqrt{\lambda }x)"

"y'=-A\sqrt{\lambda}\sin(\sqrt{\lambda }x)+B\sqrt{\lambda}\cos(\sqrt{\lambda }x)"

"y'(0)=0=>B=0"

"y'(3)=0=>\sin(3\sqrt{\lambda })=0"

"3\sqrt{\lambda }=\pi n, n=1, 2, 3,..."

"\lambda_n=\dfrac{n^2\pi ^2}{9},n=1, 2, 3,..."

The eigenvalue problem has the eigenvalue "\lambda_0=0," with associated eigenfunction "y_0=1," and infinitely many positive eigenvalues "\lambda_n=\dfrac{n^2\pi ^2}{9}," with associated eigenfunctions "y_n=\cos(\dfrac{n\pi x}{3}), n=1,2,3,..."

There are no other eigenvalues.